Find (f+g)(x),(f-g)(x),(fg)(x),and their domains.
f(x)=x+3,g(x)=x-3
(f+g)(x)=
(Simplify your answer.)

Answer by Gogonati(855)

(f+g)(x)=x+3+x-3=2x
(f-g)(x)=x+3-x+3=6
(f.g)(x)=(x+3)(x-3)=x^2-9.
Each from the functions have for domain the set of real numbers.